# Infinite-Dimensional Stochastic Differential Equations and Diffusion Dynamics of Coulomb Random Point Fields

### v3

@@ 8. Weak solutions of ISDE for the lower Dirichlet form
Assume ((A1))((A3)) and (2.60). Then, for $\mu$-a.s. $\mathsf{s}$, the sequence $\underline{\mathbf{X}}_{R}^{N_{\mathfrak{n}}}$ converges in finitely dimensional distributions to $\underline{\mathbf{X}}_{R}$ as $\mathfrak{n}\to\infty$.
@@
Let $k=\mathsf{s}(\overline{S}_{R})$. Suppose that $\mathfrak{n}$ satisfies $N_{\mathfrak{n}}\geq k$. We regard $\underline{\mathbf{X}}_{R}^{N_{\mathfrak{n}}}$ as the $\overline{S}_{R}^{k}$-valued diffusion with the reflecting boundary condition. To prove Lemma 12.1, it is sufficient to show the convergence of finitely dimensional distributions for each $k$.
Let $\mathfrak{m}_{R,\mathsf{s}}^{N_{\mathfrak{n}},k}$ and $\mathfrak{m}_{R,\mathsf{s}}^{k}$ be the $k$-density functions of $\mu_{R,\mathsf{s}}^{N_{\mathfrak{n}}}$ and $\mu_{R,\mathsf{s}}$ on $\overline{S}_{R}^{k}$, respectively. Recall that $\mathfrak{m}_{R,\mathsf{s}}^{N_{\mathfrak{n}},k}(\mathbf{x})$ and $\mathfrak{m}_{R,\mathsf{s}}^{k}(\mathbf{x})$ are symmetric in $\mathbf{x}$. Hence, according to Lemma 5.3, we deduce that, for each $l$, $\mathfrak{m}_{R,\mathsf{s}}^{N_{\mathfrak{n}},k}\wedge l$ converges to $\mathfrak{m}_{R,\mathsf{s}}^{k}\wedge l$ in $L^{2}(\overline{S}_{R}^{k},d\mathbf{x})$ as $\mathfrak{n}\to\infty$.
Let $p_{R,\mathsf{s}}^{N_{\mathfrak{n}},k}(t,\mathbf{x},\mathbf{y})\mathfrak{m}_{R,\mathsf{s}}^{N_{\mathfrak{n}},k}(\mathbf{y})d\mathbf{y}$ be the transition probability of $\underline{\mathbf{X}}_{R}^{N_{\mathfrak{n}}}$. Similarly, let $p_{R,\mathsf{s}}^{k}(t,\mathbf{x},\mathbf{y})\mathfrak{m}_{R,\mathsf{s}}^{k}(\mathbf{y})d\mathbf{y}$ be the transition probability of $\underline{\mathbf{X}}_{R}$.
Let $\mathscr{N}=\{\mathbf{x}=({x}^{1},\ldots,{x}^{k})\in\overline{S}_{R}^{k};x^{i}\neq x^{j}\text{ for }i\neq j\}$ and $\mathscr{O}=(0,\infty)\times(\overline{S}_{R}^{k}\backslash\mathscr{N})^{2}$. Then, using the classical theory of heat kernels with divergence form, we obtain that $p_{R,\mathsf{s}}^{N_{\mathfrak{n}},k}(t,\mathbf{x},\mathbf{y})$ and $p_{R,\mathsf{s}}^{k}(t,\mathbf{x},\mathbf{y})$ are locally Hölder continuous on $\mathscr{O}$ and that $p_{R,\mathsf{s}}^{N_{\mathfrak{n}},k}(t,\mathbf{x},\mathbf{y})$ converges to $p_{R,\mathsf{s}}^{k}(t,\mathbf{x},\mathbf{y})$ uniformly on each compact set $K$ in $\mathscr{O}$. Hence, we obtain the convergence of the finitely dimensional distributions of $\underline{\mathbf{X}}_{R}^{N_{\mathfrak{n}}}$ to those of $\underline{\mathbf{X}}_{R}$. ∎
@@
Assume ((A1))((A3)), (2.61), and (2.62). Then, the sequence $\underline{\mathbf{X}}_{R}^{N_{\mathfrak{n}}}$ is tight in $C([0,\infty);\mathbb{R}^{d})^{\mathbb{N}}$ as $\mathfrak{n}\to\infty$.
@@
We suppose $N_{\mathfrak{n}}\geq\mathsf{s}(\overline{S}_{R})$. We only consider the case $i\leq\mathsf{s}(\overline{S}_{R})$ because $\underline{X}_{R}^{N_{\mathfrak{n}},i}(t)$ is frozen for $i>\mathsf{s}(\overline{S}_{R})$. From (2.56), we have that, for $i\leq\mathsf{s}(\overline{S}_{R})$,
$\underline{X}_{R}^{N_{\mathfrak{n}},i}(t)-\underline{X}_{R}^{N_{\mathfrak{n}},i}(0)$ $=\int_{0}^{t}\sigma(\underline{X}_{R}^{N_{\mathfrak{n}},i}(u))dB_{u}^{i}+\int_{0}^{t}\mathfrak{b}^{N_{\mathfrak{n}}}(\underline{X}_{R}^{N_{\mathfrak{n}},i}(u),\sum_{j\neq i}^{N_{\mathfrak{n}}}\delta_{\underline{X}_{R}^{N_{\mathfrak{n}},j}(u)})du$
$+\int_{0}^{t}\frac{1}{2}\mathbf{n}_{R}\mathfrak{a}(\underline{X}_{R}^{N_{\mathfrak{n}},i}(u))dL_{R}^{N_{\mathfrak{n}},i}(u),$
$L_{R}^{N_{\mathfrak{n}},i}(t)$ $=\int_{0}^{t}1_{\partial S_{R}}(\underline{X}_{R}^{N_{\mathfrak{n}},i}(u))dL_{R}^{N_{\mathfrak{n}},i}(u).$
Note that the unlabeled dynamics $\underline{\mathsf{X}}_{R}^{N_{\mathfrak{n}}}$ are $\mu_{R,\mathsf{s}}^{N_{\mathfrak{n}}}$-reversible. Suppose that $\varphi=\mathrm{const}$. Then, from the Lyons-Zheng decomposition [5, p.284], for $1\leq i\leq\mathsf{s}(\overline{S}_{R})$,
$\underline{X}_{R}^{N_{\mathfrak{n}},i}(t)-\underline{X}_{R}^{N_{\mathfrak{n}},i}(0)=\frac{1}{2}\Big{\{}\int_{0}^{t}\sigma(\underline{X}_{R}^{N_{\mathfrak{n}},i}(v))dB_{v}^{i}-\int_{0}^{t}\sigma(\underline{X}_{R}^{N_{\mathfrak{n}},i}(v))dB_{v}^{i}\circ r_{t}\Big{\}}.$ (12.1)
From (12.1) and (2.39), and the martingale inequality, we obtain a uniform bound
[…]
Hence, $\underline{\mathbf{X}}_{R}^{N_{\mathfrak{n}}}$ is tight in $C([0,\infty);\mathbb{R}^{d})^{\mathbb{N}}$ as $\mathfrak{n}\to\infty$, where $c_{12.1}$ is a constant independent of $\mathfrak{n}$. For $\varphi$ as in (2.61), the tightness of $\underline{\mathbf{X}}_{R}^{N_{\mathfrak{n}}}$ follows from (2.62) and
[…]
We have thus completed the proof. ∎
Proof of Theorem 2.5.
From Lemma 12.1 and Lemma 12.2, we have (2.64). From (8.11), we have (2.65). We have thus completed the proof. ∎

### v4

@@ 12.1. Finite-domain approximation
i The sequence $\underline{\mathbf{X}}_{R}^{N}$ $(N,R\in\mathbb{N})$ is tight in $W(\mathbb{R}^{d})^{\mathbb{N}}$.
ii The sequence $\mathbf{X}^{N,[m]}$ $(N\in\mathbb{N})$ is tight in $W((\mathbb{R}^{d})^{m}\times\mathsf{S})$ for $m\in\{0\}\cup\mathbb{N}$.
@@
We suppose $N\geq\mathsf{s}(\overline{S}_{R})$. We only consider the case $i\leq\mathsf{s}(\overline{S}_{R})$ because $\underline{X}_{R}^{N,i}(t)$ is frozen for $i>\mathsf{s}(\overline{S}_{R})$.
The unlabeled dynamics $\underline{\mathsf{X}}_{R}^{N}$ are $\mu_{R,\mathsf{s}}^{N}$-reversible. Suppose that $\varphi=\mathrm{const}$. Then from the Lyons-Zheng decomposition [5, p.284],
$\underline{X}_{R}^{N,i}(t)-\underline{X}_{R}^{N,i}(0)=\frac{1}{2}\Big\{\int_{0}^{t}\sigma(\underline{X}_{R}^{N,i}(v))dB_{v}^{i}-\int_{0}^{t}\sigma(\underline{X}_{R}^{N,i}(v))dB_{v}^{i}\circ r_{t}\Big\}.$
From this, (2.31), (2.44)–(2.45), and the martingale inequality, we obtain
[…]
where $c_{12.1}$ is a constant independent of $N,R\in\mathbb{N}$. By the Lyons-Zheng decomposition, $\mathbf{X}^{N}$ is tight in $W(\mathbb{R}^{d})^{\mathbb{N}}$. This implies i.
Moreover, by the Lyons-Zheng decomposition, any limit point satisfies ((NBJ)). From this and Lemma 13.3, we obtain ii. ∎
Recall that, for $\mathsf{y}\in\mathsf{S}$, $\mathsf{y}^{N_{\mathfrak{n}}}=(\mathfrak{l}^{1}(\mathsf{y}),\ldots,\mathfrak{l}^{N_{\mathfrak{n}}}(\mathsf{y}))$, where $\mathfrak{l}=(\mathfrak{l}^{i})_{i}$ is the label as in (2.22). Thus, $\mathsf{y}^{N_{\mathfrak{n}}}$ is a function of $\mathsf{y}$.
@@
For $\mu$-a.s. $\mathsf{y}$, the following hold:
i The Dirichlet form $(\mathscr{E}_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}},[m]},\underline{\mathscr{D}}_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}},[m]})$ on $L^{2}(\mu_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}},[m]})$ converges to the Dirichlet form $(\mathscr{E}_{R,\mathsf{y}}^{[m]},\underline{\mathscr{D}}_{R,\mathsf{y}}^{[m]})$ on $L^{2}(\mu_{R,\mathsf{y}}^{[m]})$ in the strong resolvent sense after identifying $L^{2}(\mu_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}},[m]})$ and $L^{2}(\mu_{R,\mathsf{y}}^{[m]})$ with their images in $L^{2}(\Lambda_{R}^{[m]})$ under the canonical isometric embeddings induced by the Radon–Nikodym derivatives.
ii The Dirichlet form $(\mathscr{E}_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}},[m]},\overline{\mathscr{D}}_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}},[m]})$ on $L^{2}(\mu_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}},[m]})$ converges to the Dirichlet form $(\mathscr{E}_{R,\mathsf{y}}^{[m]},\overline{\mathscr{D}}_{R,\mathsf{y}}^{[m]})$ on $L^{2}(\mu_{R,\mathsf{y}}^{[m]})$ in the strong resolvent sense on $L^{2}(\Lambda_{R}^{[m]})$ under the identification in i.
@@
Let $\mathfrak{m}_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}},k}$ and $\mathfrak{m}_{R,\mathsf{y}}^{k}$ be the $k$-density functions of $\mu_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}}}$ and $\mu_{R,\mathsf{y}}$, respectively, on $\overline{S}_{R}^{k}$. By (5.31), $\mathfrak{m}_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}},k}$ converges to $\mathfrak{m}_{R,\mathsf{y}}^{k}$ in $C(\overline{S}_{R}^{k})$.
Let $p_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}},k}(t,\mathbf{x},\mathbf{y})\mathfrak{m}_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}},k}(\mathbf{y})\,d\mathbf{y}$ and $p_{R,\mathsf{y}}^{k}(t,\mathbf{x},\mathbf{y})\mathfrak{m}_{R,\mathsf{y}}^{k}(\mathbf{y})\,d\mathbf{y}$ be the transition probabilities of $\underline{\mathbf{X}}_{R}^{N_{\mathfrak{n}}}$ and $\underline{\mathbf{X}}_{R}$ on $\overline{S}_{R}^{k}$, respectively. For each $R\in\mathbb{N}$ and $k\in\mathbb{N}$, these transition densities exist by the classical theory of heat kernels in divergence form. We set
$S_{\neq}$ $=\bigl\{\mathbf{x}=(x^{i})_{i=1}^{k}\in\overline{S}_{R}^{k}\,;\,x^{i}\neq x^{j}\text{ for }i\neq j\bigr\},\quad\mathscr{O}=(0,\infty)\times S_{\neq}^{2}.$
By the classical theory of heat kernels in divergence form, $p_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}},k}$ and $p_{R,\mathsf{y}}^{k}$ are locally Hölder continuous on $\mathscr{O}$, and $p_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}},k}$ converges to $p_{R,\mathsf{y}}^{k}$ uniformly on each compact subset of $\mathscr{O}$. Moreover, $\mathfrak{m}_{R,\mathsf{y}^{N_{\mathfrak{n}}}}^{N_{\mathfrak{n}},k}$ converges to $\mathfrak{m}_{R,\mathsf{y}}^{k}$ in $C(\overline{S}_{R}^{k})$, as noted at the beginning of the proof. Consequently, the corresponding $\alpha$-resolvent densities converge, and i follows.
The Dirichlet form in ii is associated with a different boundary condition. Nevertheless, its heat kernel converges locally uniformly in the same manner as that in i. This implies the assertion of ii. ∎
Let $(\mathscr{E}_{R}^{N_{\mathfrak{n}},[m]},\underline{\mathscr{D}}_{R\star}^{N_{\mathfrak{n}},[m]})$ and $(\mathscr{E}_{R}^{N_{\mathfrak{n}},[m]},\overline{\mathscr{D}}_{R\star}^{N_{\mathfrak{n}},[m]})$ be the superpositions, with respect to $\mathsf{y}$ under $\mu$, of $(\mathscr{E}_{R,\mathsf{y}}^{N_{\mathfrak{n}},[m]},\underline{\mathscr{D}}_{R,\mathsf{y}}^{N_{\mathfrak{n}},[m]})$ and $(\mathscr{E}_{R,\mathsf{y}}^{N_{\mathfrak{n}},[m]},\overline{\mathscr{D}}_{R,\mathsf{y}}^{N_{\mathfrak{n}},[m]})$, respectively.
Lemma 12.3.
i The Dirichlet form $(\mathscr{E}_{R}^{N_{\mathfrak{n}},[m]},\underline{\mathscr{D}}_{R\star}^{N_{\mathfrak{n}},[m]})$ on $L^{2}(\mu_{R}^{N_{\mathfrak{n}},[m]})$ converges to the Dirichlet form $(\mathscr{E}_{R}^{[m]},\underline{\mathscr{D}}_{R\star}^{[m]})$ on $L^{2}(\mu_{R}^{[m]})$ in the strong resolvent sense on $L^{2}(\Lambda_{R}^{[m]})$ after identifying $L^{2}(\mu_{R}^{N_{\mathfrak{n}},[m]})$ and $L^{2}(\mu_{R}^{[m]})$ with their images in $L^{2}(\Lambda_{R}^{[m]})$ under the canonical isometric embeddings induced by the Radon–Nikodym derivatives.
ii The Dirichlet form $(\mathscr{E}_{R}^{N_{\mathfrak{n}},[m]},\overline{\mathscr{D}}_{R\star}^{N_{\mathfrak{n}},[m]})$ on $L^{2}(\mu_{R}^{N_{\mathfrak{n}},[m]})$ converges to the Dirichlet form $(\mathscr{E}_{R}^{[m]},\overline{\mathscr{D}}_{R\star}^{[m]})$ on $L^{2}(\mu_{R}^{[m]})$ in the strong resolvent sense on $L^{2}(\Lambda_{R}^{[m]})$ under the identification in i.
Proof.
By the definition of superposition together with Fubini’s theorem and Jensen’s inequality, Lemma 12.3 follows from Lemma 12.2. ∎
We define the particles with indices greater than $N_{n}$ to remain fixed for all time. With this convention, the dynamics of the first $N_{n}$ particles can be regarded as an infinite-particle system. Under this convention, all solutions start from the same initial condition $\mathfrak{l}(\mathsf{s})$ almost surely.
Theorem 12.1 (Finite-domain approximation).
Assume ((A1))((A3)) and (2.44)–(2.45). Then
$\lim_{\mathfrak{n}\to\infty}\underline{\mathbf{X}}_{R}^{N_{\mathfrak{n}}}$ $=\underline{\mathbf{X}}_{R}\quad\text{ in law in }W(\mathbb{R}^{d})^{\mathbb{N}},$ (12.3)
$\lim_{R\to\infty}\underline{\mathbf{X}}_{R}$ $=\mathbf{X}\quad\text{ in law in }W(\mathbb{R}^{d})^{\mathbb{N}}.$ (12.4)
Proof.
Since $\underline{\mathbf{X}}_{R}$ is a solution of SDE (8.1)–(8.2), it follows that $\underline{\mathbf{X}}_{R}^{[m]}$ is associated with $(\mathscr{E}_{R}^{[m]},\underline{\mathscr{D}}_{R\star}^{[m]})$ on $L^{2}(\mu_{R}^{[m]})$. Hence, by Lemma 12.1 i and Lemma 12.3 i, we obtain (12.3).
By Theorem 8.1 iii, we obtain (12.4), which completes the proof. ∎
12.2. Finite-particle approximation: Proof of Theorem 2.5
Proof of Theorem 2.5. Let $\mathbf{X}^{N}$ be the solution of SDE (2.42) with (2.44)–(2.45). For each $m\in\{0\}\cup\mathbb{N}$ with $m\leq N$, let
$\mathscr{E}^{N,[m]}(f,g)=\int_{(\mathbb{R}^{d})^{m}\times\mathsf{S}}\mathbb{D}_{\mathfrak{a}}^{[m]}[f,g]d\mu^{N,[m]}.$
Then $(\mathscr{E}^{N,[m]},\mathscr{D}_{\circ\mathrm{b}}^{[m]})$ is closable on $L^{2}(\mu^{N,[m]})$. Let $(\mathscr{E}^{N,[m]},\mathscr{D}^{N,[m]})$ be its closure. The diffusion $\mathbf{X}^{N,[m]}$ is properly associated with $(\mathscr{E}^{N,[m]},\mathscr{D}^{N,[m]})$ on $L^{2}(\mu^{N,[m]})$. By the Lyons–Zheng decomposition and (2.44)–(2.45), $\mathbf{X}^{N}$ and $\mathbf{X}^{N,[m]}$, $N\in\mathbb{N}$, are tight in $W(\mathbb{R}^{d})^{\mathbb{N}}$ and $W((\mathbb{R}^{d})^{m}\times\mathsf{S})$, respectively.
Hence, there exist subsequences of $\mathbf{X}^{N_{\mathfrak{n}}}$ and $\mathbf{X}^{N_{\mathfrak{n}},[m]}$, still denoted by the same symbols, and limits $\widetilde{\mathbf{X}}$ and $\widetilde{\mathbf{X}}^{[m]}$ such that $\widetilde{\mathbf{X}}^{[m]}$ is the $m$-labeled process of $\widetilde{\mathbf{X}}$. Furthermore, for any nonnegative $f,g\in\mathscr{D}_{\circ\mathrm{b}}^{[m]}$ and $\alpha>0$,
$\lim_{\mathfrak{n}\to\infty}\int_{(\mathbb{R}^{d})^{m}\times\mathsf{S}}fR^{\alpha,N_{\mathfrak{n}},[m]}g\,d(\mu^{[m]})^{2}=\int_{(\mathbb{R}^{d})^{m}\times\mathsf{S}}f\widetilde{R}^{\alpha,[m]}g\,d(\mu^{[m]})^{2}.$ (12.5)
Here $R^{\alpha,N_{\mathfrak{n}},[m]}$ denotes the $\alpha$-resolvent of $(\mathscr{E}^{N,[m]},\mathscr{D}^{N,[m]})$ on $L^{2}(\mu^{N,[m]})$ and
$\widetilde{R}^{\alpha,[m]}g(\mathbf{x},\mathsf{s})=E_{(\mathbf{x},\mathsf{s})}[\int_{0}^{\infty}\mathrm{e}^{-\alpha t}g(\widetilde{\mathbf{X}}_{t}^{[m]})dt].$
At this stage, it is not known whether $\widetilde{\mathbf{X}}^{[m]}$ is a Markov process. We now prove that this is indeed the case.
Let $(\mathscr{E}_{R}^{N_{\mathfrak{n}},[m]},\underline{\mathscr{D}}_{R\star}^{N_{\mathfrak{n}},[m]})$ and $(\mathscr{E}_{R}^{N_{\mathfrak{n}},[m]},\overline{\mathscr{D}}_{R\star}^{N_{\mathfrak{n}},[m]})$ be as in Lemma 12.3.
For symmetric bilinear forms $A$ and $B$, $A\succ B$ means that $A$ is an extension of $B$ in the sense of (13.1). Then it is clear that
$(\mathscr{E}_{R}^{N_{\mathfrak{n}},[m]},\underline{\mathscr{D}}_{R\star}^{N_{\mathfrak{n}},[m]})\succ(\mathscr{E}^{N_{\mathfrak{n}},[m]},\mathscr{D}^{N_{\mathfrak{n}},[m]})\succ(\mathscr{E}_{R}^{N_{\mathfrak{n}},[m]},\overline{\mathscr{D}}_{R\star}^{N_{\mathfrak{n}},[m]}).$
Hence,
$(\mathscr{E}_{R}^{N_{\mathfrak{n}},[m]},\underline{\mathscr{D}}_{R\star}^{N_{\mathfrak{n}},[m]})\leq(\mathscr{E}^{N_{\mathfrak{n}},[m]},\mathscr{D}^{N_{\mathfrak{n}},[m]})\leq(\mathscr{E}_{R}^{N_{\mathfrak{n}},[m]},\overline{\mathscr{D}}_{R\star}^{N_{\mathfrak{n}},[m]}).$ (12.6)
By Lemma 12.3 i, $(\mathscr{E}_{R}^{N_{\mathfrak{n}},[m]},\underline{\mathscr{D}}_{R\star}^{N_{\mathfrak{n}},[m]})$ converges to $(\mathscr{E}_{R}^{[m]},\underline{\mathscr{D}}_{R\star}^{[m]})$ in the strong resolvent sense as $\mathfrak{n}\to\infty$. By Lemma 13.5 i, $(\mathscr{E}_{R}^{[m]},\underline{\mathscr{D}}_{R\star}^{[m]})$ converges to the increasing limit $(\mathscr{E}^{[m]},\mathscr{D}_{\mathrm{lwr}\star}^{[m]})$ in the strong resolvent sense as $R\to\infty$. Thus, the left closed form in (12.6) converges to $(\mathscr{E}^{[m]},\mathscr{D}_{\mathrm{lwr}\star}^{[m]})$ in the strong resolvent sense in the two-step limits.
Let $f\in\overline{\mathscr{D}}_{R}^{N_{\mathfrak{n}},[m]}$. Then $f$ is $\mathscr{B}((S_{R})^{m})\times\sigma[\pi_{S_{R}}]$-measurable. Moreover, $\mu_{R}^{N_{\mathfrak{n}},[m]}$ is absolutely continuous with respect to $\mu_{R}^{[m]}$. Hence, we regard $(\mathscr{E}_{R}^{N_{\mathfrak{n}},[m]},\overline{\mathscr{D}}_{R\star}^{N_{\mathfrak{n}},[m]})$ as a closed form on $L^{2}(\mu_{R}^{[m]})$.
By Lemma 12.3 ii, $(\mathscr{E}_{R}^{N_{\mathfrak{n}},[m]},\overline{\mathscr{D}}_{R\star}^{N_{\mathfrak{n}},[m]})$ converges to $(\mathscr{E}_{R}^{[m]},\overline{\mathscr{D}}_{R\star}^{[m]})$ in the strong resolvent sense. Moreover, by Lemma 13.5 ii, $(\mathscr{E}_{R}^{[m]},\overline{\mathscr{D}}_{R\star}^{[m]})$ converges to the decreasing limit $(\mathscr{E}^{[m]},\mathscr{D}_{\mathrm{upr}\star}^{[m]})$ in the strong resolvent sense on $L^{2}(\mu^{[m]})$. Thus, the right closed form in (12.6) converges to $(\mathscr{E}^{[m]},\mathscr{D}_{\mathrm{upr}\star}^{[m]})$ in the strong resolvent sense in the two-step limits.
Hence, any limit point of $(\mathscr{E}^{N_{\mathfrak{n}},[m]},\mathscr{D}^{N_{\mathfrak{n}},[m]})$ lies between $(\mathscr{E}^{[m]},\mathscr{D}_{\mathrm{lwr}\star}^{[m]})$ and $(\mathscr{E}^{[m]},\mathscr{D}_{\mathrm{upr}\star}^{[m]})$ in the sense that, for any such nonnegative $f,g\in\mathscr{D}_{\circ\mathrm{b}}$, the quantity in the right-hand side of (12.5) is between $\int_{\mathsf{S}}f\underline{R}_{\star,\infty}^{\alpha,[m]}gd\mu^{[m]}$ and $\int_{\mathsf{S}}f\overline{R}_{\star,\infty}^{\alpha,[m]}gd\mu^{[m]}$. Here $\underline{R}_{\star,\infty}^{\alpha,[m]}$ and $\overline{R}_{\star,\infty}^{\alpha,[m]}$ are the $\alpha$-resolvent of $(\mathscr{E}^{[m]},\mathscr{D}_{\mathrm{lwr}\star}^{[m]})$ and $(\mathscr{E}^{[m]},\mathscr{D}_{\mathrm{upr}\star}^{[m]})$ , respectively.
By Theorem 11.1, these two Dirichlet forms coincide. We write their common form as $(\mathscr{E}^{[m]},\mathscr{D}_{\star}^{[m]})$, and denote its $\alpha$-resolvent on $L^{2}(\mu^{[m]})$ by $R_{\star}^{\alpha,[m]}$. Therefore, $(\mathscr{E}^{N_{\mathfrak{n}},[m]},\mathscr{D}^{N_{\mathfrak{n}},[m]})$ converges to $(\mathscr{E}^{[m]},\mathscr{D}_{\star}^{[m]})$ in the following sense. For any nonnegative $f,g\in\mathscr{D}_{\circ\mathrm{b}}$,
$\lim_{\mathfrak{n}\to\infty}\int_{(\mathbb{R}^{d})^{m}\times\mathsf{S}}fR^{\alpha,N_{\mathfrak{n}},[m]}g\,d(\mu^{[m]})^{2}=\int_{(\mathbb{R}^{d})^{m}\times\mathsf{S}}fR_{\star}^{\alpha,[m]}g\,d\mu^{[m]}.$
Hence $\widetilde{R}^{\alpha,[m]}=R_{\star}^{\alpha,[m]}$. This implies that $\widetilde{\mathbf{X}}^{[m]}$ is a Markov process with resolvents $\widetilde{R}^{\alpha,[m]}$, $\alpha>0$, and that $\widetilde{\mathbf{X}}^{[m]}=\mathbf{X}^{[m]}$ in law. Hence
$\lim_{\mathfrak{n}\to\infty}\mathbf{X}^{N_{\mathfrak{n}},[m]}=\mathbf{X}^{[m]}\quad\text{in finite-dimensional distributions}.$ (12.7)
This, together with tightness in $W(\mathbb{R}^{d})^{\mathbb{N}}$, completes the proof. ∎
Remark 12.2.
The main difficulty in proving Theorem 2.5 is that the state space of the $N$-particle system depends on $N$. In [16], using a general theorem on the convergence of Dirichlet forms in distinct state spaces [19], a result similar to Theorem 2.5 was proved. In that work, the $N$-particle dynamics is considered in domains $S_{R_{N}}$ with reflecting boundary conditions such that $R_{N}\to\infty$. However, this does not yield convergence of the particle systems in the whole space and is therefore unsatisfactory.
The key idea in the proof of Theorem 2.5 is the identification of a relation connecting the $N$-particle Dirichlet form with finite-domain Dirichlet forms.